# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v4

@@ B.6 Additional PEFT baselines for MAE
As a continuation of table 1, we include PEFT methods used on MAE weights, which generally underperform compared with DinoV2. For completeness, these results are in table 16.
@@
Table 16: MAE+PEFT results on fMoW-RGB validation split (table 1, contd.). “Pre-train Params" and “Fine-tune Params" refer to trainable parameters of the ViT encoder required on the new domain (satellite images).
@@ B.7 Results on UDA Benchmarks
As discussed in section A.2, while ExPLoRA is not a traditional unsupervised domain adaptation (UDA) method, it can serve as an initialization for ViT-based UDA approaches. We demonstrate this compatibility below.
@@
Table 17 shows ExPLoRA’s effectiveness when combined with TVT (Yang et al., 2023) and SSRT (Sun et al., 2022), two state-of-the-art UDA methods. Using ExPLoRA D-[12]-$r64$ (DinoV2-initialized ViT-B with last layer unfrozen and LoRA-r64 elsewhere) pre-trained on both synthetic and real domains, we outperform traditional ImageNet-21k (Deng et al., 2009) initialization by 1-3% while achieving more balanced per-class accuracy. Most notably, when using SSRT, we achieve a new SoTA accuracy of 90.9%, surpassing both ImageNet and DinoV2 initialization by substantial margins ($\uparrow$2.1and $\uparrow$1.0respectively). These results are particularly significant as they show ExPLoRA’s unsupervised initialization can outperform methods that rely on supervised ImageNet-21k pre-training. The benefits of ExPLoRA initialization are also clear with TVT, where performance rises to 88.2% to match recent SoTA methods (Xu et al., 2021; Zhu et al., 2023)– a marked improvement over both its original results ($\uparrow$2.1%) and DinoV2 initialization ($\uparrow$4.1%). These results demonstrate that ExPLoRA serves as a powerful initialization strategy for UDA, effectively bridging domain gaps while remaining computationally efficient.
@@ B.8 Attention Map Visualizations
To aid our analysis in section 7, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table 3), the second and third rows of table 3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig. 9 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2024).
@@
The visualizations in fig. 9 further support the analysis in section 7. For the Dino models, the attention scores of block 9-10 are diffuse and spread around the central object of the image, with quite a few border pixels highlighted. Conversely, the attention scores of the final layers are concentrated more towards the central object. These visualizations further suggest that the middle layers focus on capturing local properties of the images such as texture, while the final layers capture global semantic information such as object-ness. Interestingly, the initial blocks for the Dino models display sparse attention patterns with spikes on seemingly random patches. This might suggest a form of caching to aid the computation of deeper layers that will extract local or global information.
@@ C.1 Pre-Training
We use the ViT-Large architecture for all experiments. Since raw image sizes vary, the shorter image size is resized to $224$ while preserving aspect ratio, and then a center crop is taken to yield images of size $3\times 224\times 224$, representing the channels, height, and width. All pre-training is done on a single NVIDA-RTX 6000 Ada GPU, or 4 NVIDIA-RTX A4000 GPUs on an academic GPU cluster.
@@ C.2 PEFT Fine-Tuning
For data augmentations, we only use the drop-path augmentation (Larsson et al., 2016) at a rate of 0.2, with no dropout, mixup, or cutmix. We note that the original LoRA configuration outperforms other PEFT techniques when paired with the drop-path regularization technique. For example, we find that BOFT does not pair well with drop-path, instead performing most effectively with a custom multiplicative dropout technique (Liu et al., 2024). We include the result with the best hyperparameter configuration for each row in table 1.
@@ C.4 Multi-Spectral Images
We use the group-channel ViT-L architecture introduced in (Cong et al., 2022). We don’t use DinoV2 since there is no such architecture for DinoV2 pre-training. Input images are $13\times 98\times 98$, representing 13 multi-spectral bands. We follow the configuration in Cong et al. (2022) of dropping bands B1, B9, B10, and use the same grouping strategy. When loading MAE weights to the ViT-L encoder, the patch embeddings do not match and so the patch embedding and group channel encodings are trained from scratch. All other configuration details are the same as for M-ExPLoRA in section C.1, except that we use a base learning rate of $4.5\times 10^{-4}$ for pre-training and train for 50 epochs (given the larger dataset size) on 4 NVIDIA RTX A4000 GPUs for 80 hours.
@@ C.5 Dataset Information
We include detailed information about the datasets used in our experiments. Table 18 shows the number of samples in the train and validation splits for all datasets used in this work.
Hyperparameter and training configuration details are the same as in section C.1 if the images are RGB, and the same as in section C.4 if the images have more channels or are temporal.
Since each dataset contains a varying number of training images, the number of ExPLoRA pre-training iterations should be adjusted accordingly. We include the recommended number of ExPLoRA pre-training iterations in table 18. Note, however, that this number can be varied depending on other hyperparameters such as batch size, learning rate, LoRA rank, and number of unfrozen blocks.
@@
The licenses for all datasets are included in the footnotes: fMoW111fMoW license: https://github.com/fMoW/dataset/raw/master/LICENSE, Sentinel-$2$Sentinel-2 license: https://scihub.copernicus.eu/twiki/pub/SciHubWebPortal/TermsConditions/Sentinel_Data_Terms_and_Conditions.pdf, EuroSAT [^3] EuroSAT license: https://creativecommons.org/licenses/by/4.0/, SpaceNet [^4] SpaceNet v1 license: http://creativecommons.org/licenses/by-sa/4.0/, Camelyon17555Camelyon17 license:https://creativecommons.org/publicdomain/zero/1.0/, iWildCam666iWildCam license:https://cdla.dev/permissive-1-0/, GlobalWheat [^7] GlobalWheat license:https://opensource.org/licenses/MIT.
@@ Appendix D Environmental Impact
Following (Cong et al., 2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table 19888Note: while the SatMAE paper reported 768 GPU hours for pre-training on fMoW-RGB and 576 GPU hours on fMoW-Sentinel, our measurements in Tables 1-5 of the main text have a revised number to ensure consistency of hardware platforms across comparisons.
@@
Since we initialize with pre-trained weights on natural image domains, ExPLoRA is much less environmentally impactful while achieving similar or higher levels of performance. We achieve a 4x-8x reduction in total carbon emitted for each of the large pre-training satellite image datasets considered in table 19.

### v5

@@ B.6 Additional PEFT baselines for MAE
As a continuation of table˜1, we include PEFT methods used on MAE weights, which generally underperform compared with DinoV2. For completeness, these results are in table˜16.
@@
Table 16: MAE+PEFT results on fMoW-RGB validation split (table˜1, contd.). “Pre-train Params" and “Fine-tune Params" refer to trainable parameters of the ViT encoder required on the new domain (satellite images).
@@ B.7 Results on UDA Benchmarks
As discussed in section˜A.2, while ExPLoRA is not a traditional unsupervised domain adaptation (UDA) method, it can serve as an initialization for ViT-based UDA approaches. We demonstrate this compatibility below.
@@
Table˜17 shows ExPLoRA’s effectiveness when combined with TVT (Yang et al., 2023) and SSRT (Sun et al., 2022), two state-of-the-art UDA methods. Using ExPLoRA D-[12]-$r64$ (DinoV2-initialized ViT-B with last layer unfrozen and LoRA-r64 elsewhere) pre-trained on both synthetic and real domains, we outperform traditional ImageNet-21k (Deng et al., 2009) initialization by 1-3% while achieving more balanced per-class accuracy. Most notably, when using SSRT, we achieve a new SoTA accuracy of 90.9%, surpassing both ImageNet and DinoV2 initialization by substantial margins ($\uparrow$2.1% and $\uparrow$1.0% respectively). These results are particularly significant as they show ExPLoRA’s unsupervised initialization can outperform methods that rely on supervised ImageNet-21k pre-training. The benefits of ExPLoRA initialization are also clear with TVT, where performance rises to 88.2% to match recent SoTA methods (Xu et al., 2021; Zhu et al., 2023)– a marked improvement over both its original results ($\uparrow$2.1%) and DinoV2 initialization ($\uparrow$4.1%). These results demonstrate that ExPLoRA serves as a powerful initialization strategy for UDA, effectively bridging domain gaps while remaining computationally efficient.
@@ B.8 Attention Map Visualizations
To aid our analysis in section˜7, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table˜3), the second and third rows of table˜3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig.˜9 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2024).
@@
The visualizations in fig.˜9 further support the analysis in section˜7. For the Dino models, the attention scores of block 9-10 are diffuse and spread around the central object of the image, with quite a few border pixels highlighted. Conversely, the attention scores of the final layers are concentrated more towards the central object. These visualizations further suggest that the middle layers focus on capturing local properties of the images such as texture, while the final layers capture global semantic information such as object-ness. Interestingly, the initial blocks for the Dino models display sparse attention patterns with spikes on seemingly random patches. This might suggest a form of caching to aid the computation of deeper layers that will extract local or global information.
@@ C.1 Pre-Training
We use the ViT-Large architecture for all experiments. Since raw image sizes vary, the shorter image size is resized to $224$ while preserving aspect ratio, and then a center crop is taken to yield images of size $3\times 224\times 224$, representing the channels, height, and width. All pre-training is done on a single NVIDIA-RTX 6000 Ada GPU, or 4 NVIDIA-RTX A4000 GPUs on an academic GPU cluster.
@@ C.2 PEFT Fine-Tuning
For data augmentations, we only use the drop-path augmentation (Larsson et al., 2016) at a rate of 0.2, with no dropout, mixup, or cutmix. We note that the original LoRA configuration outperforms other PEFT techniques when paired with the drop-path regularization technique. For example, we find that BOFT does not pair well with drop-path, instead performing most effectively with a custom multiplicative dropout technique (Liu et al., 2024). We include the result with the best hyperparameter configuration for each row in table˜1.
@@ C.4 Multi-Spectral Images
We use the group-channel ViT-L architecture introduced in (Cong et al., 2022). We don’t use DinoV2 since there is no such architecture for DinoV2 pre-training. Input images are $13\times 98\times 98$, representing 13 multi-spectral bands. We follow the configuration in Cong et al. (2022) of dropping bands B1, B9, B10, and use the same grouping strategy. When loading MAE weights to the ViT-L encoder, the patch embeddings do not match and so the patch embedding and group channel encodings are trained from scratch. All other configuration details are the same as for M-ExPLoRA in section˜C.1, except that we use a base learning rate of $4.5\times 10^{-4}$ for pre-training and train for 50 epochs (given the larger dataset size) on 4 NVIDIA RTX A4000 GPUs for 80 hours.
@@ C.5 Dataset Information
We include detailed information about the datasets used in our experiments. Table˜18 shows the number of samples in the train and validation splits for all datasets used in this work.
Hyperparameter and training configuration details are the same as in section˜C.1 if the images are RGB, and the same as in section˜C.4 if the images have more channels or are temporal.
Since each dataset contains a varying number of training images, the number of ExPLoRA pre-training iterations should be adjusted accordingly. We include the recommended number of ExPLoRA pre-training iterations in table˜18. Note, however, that this number can be varied depending on other hyperparameters such as batch size, learning rate, LoRA rank, and number of unfrozen blocks.
@@
The licenses for all datasets are included in the footnotes: fMoW111fMoW license: https://github.com/fMoW/dataset/raw/master/LICENSE, Sentinel-2222Sentinel-2 license: https://scihub.copernicus.eu/twiki/pub/SciHubWebPortal/TermsConditions/Sentinel_Data_Terms_and_Conditions.pdf, EuroSAT [^3] EuroSAT license: https://creativecommons.org/licenses/by/4.0/, SpaceNet [^4] SpaceNet v1 license: http://creativecommons.org/licenses/by-sa/4.0/, Camelyon17555Camelyon17 license:https://creativecommons.org/publicdomain/zero/1.0/, iWildCam666iWildCam license:https://cdla.dev/permissive-1-0/, GlobalWheat [^7] GlobalWheat license:https://opensource.org/licenses/MIT.
@@ Appendix D Environmental Impact
Following (Cong et al., 2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table˜19888Note: while the SatMAE paper reported 768 GPU hours for pre-training on fMoW-RGB and 576 GPU hours on fMoW-Sentinel, our measurements in Tables 1-5 of the main text have a revised number to ensure consistency of hardware platforms across comparisons.
@@
Since we initialize with pre-trained weights on natural image domains, ExPLoRA is much less environmentally impactful while achieving similar or higher levels of performance. We achieve a 4x-8x reduction in total carbon emitted for each of the large pre-training satellite image datasets considered in table˜19.